# Orbital magnetoelectric coupling of three dimensional Chern insulators

### v1

@@ Abstract
As shown in Table 5, the quantization rule remains good statistically not only for Type-I disorder preserving layer translation symmetry but also Type-II disorder breaking it. This is because the amplitude of the disorder does not close the gap so that layer Chern number can still be well defined.
@@
Two subtleties worth mentioning in the thermodynamic limit that $\alpha_{zz}^{(l)}$ is only defined modulo $N_{\rm sub}Ce^{2}/h$ and only the difference in $\alpha_{zz}^{(l)}$ is physically measurable. Both observations are rooted in the indeterminacy of electric polarization that the choice of unit cell is arbitrary. Here $N_{\rm sub}$ is the number of inequivalent layers within the smallest unit cell. Nevertheless, the difference between $\alpha_{zz}^{(l)}$ of two neighboring layers is measurable and is a topological quantity. The ambiguity in the branch choice of $\alpha_{zz}^{(l)}$ is either fixed by the symmetry constraint on the total bulk response $\sum_{l}\alpha^{(l)}$ or by boundary condition at the surface.
In summary, we propose a diagonal layer-resolved OME response in 3D Chern insulator, which have not been studied before. Most saliently, this OME response is quantized or half-quantized to $e^{2}/h$ in the presence of inversion or mirror symmetry. Furthermore, even lack of symmetries, the gradient of this OME coupling is still exactly quantized in unit of $e^{2}/h$ given by Eq. (5). We demonstrate the robustness of the quantization rule against disorder through numerical calculations and topological argument using the Streda formula. Such considerable and quantized difference between layer-resolved OME $\alpha_{zz}^{(l)}$ is in principle measurable in experiments. For example, in a slab system, applying an out-of-plane electric field would induce a quantized change in orbital magnetization at the top and bottom layers, which may be probed by SQUID-on-tip microscope [37, 38].
@@ Acknowledgements.
This work is supported by the the National Natural Science Foundation of China (Grant No. 12174257), National Key R & D Program of China (Grant No. 2020YFA0309601).
Supplemental Materials of “Orbital magnetoelectric coupling of 3D Chern Insulator”
@@ I.2 Definition of layer-resolved orbital magnetization in the thermodynamic limit
$-\frac{e}{2\hbar}{\rm Im}\big{\{}\epsilon_{z\alpha\beta}\sum_{s\,l}\sum_{\textbf{k}}\sum_{n\in{\rm occ}}\sum_{m,m^{\prime}\in{\rm emp}}C_{sl,m}(\textbf{k})C_{sl,m^{\prime}}^{*}(\textbf{k})E_{n\textbf{k}}\langle u_{m\textbf{k}}|\partial_{k_{\alpha}}u_{n\textbf{k}}\rangle_{\Omega_{0}}\langle\partial_{k_{\beta}}u_{n\textbf{k}}|u_{m^{\prime}\textbf{k}}\rangle_{\Omega_{0}}\big{\}}$
@@ I.3 Layer Chern number
In this section, we present the numerical results that $M_{z}^{(l)}$ varies linearly with weak electric field for both types of stacking, independent of number of layers, as clearly shown in Fig. 1. Here, we calculate $M_{z}^{(l)}$ using the definition Eq. (23).

### v2

@@ Abstract
In the presence of disorder, as long as the bulk gap is not closed by disorder such that layer Chern number remains well defined and quantized, the change of layer charge density induced by magnetic field would still remain quantized as dictated by Streda formula. Thus, the quantization rule of Eq. (4) is still valid. More numerical evidence of the robustness of Eq. (4) in the presence of disorder is given in Table  5. In practice, we need to calculate $\alpha_{zz}^{(l)}$ numerically by finite difference in the open boundary conditions. Disorder is modelled by random on-site potential. Two types of disorder are introduced. The first type is that on-site energy only fluctuates in-plane repeating itself among all the layers. The second type breaks the translation symmetry such that on-site energy fluctuates among all the sites of all layers. We consider 15 disorder configurations and then analyze the results statistically. Here we set $t_{3}=0.2$, $t_{4}=0$, and a strong disorder with amplitude $0.4t_{3}$. As shown in Table 5, the quantization rule remains good statistically not only for Type-I disorder preserving layer translation symmetry but also Type-II disorder breaking it. This is because the amplitude of the disorder does not close the gap so that layer Chern number can still be well defined.
@@
It worth noting that $\alpha_{zz}^{(l)}$ is only defined modulo $N_{\rm sub}Ce^{2}/h$, which is rooted in the indeterminacy of electric polarization that the choice of unit cell is arbitrary. Here $N_{\rm sub}$ is the number of inequivalent layers within the smallest unit cell. Nevertheless, the difference between $\alpha_{zz}^{(l)}$ of two neighboring layers is physically measurable and is a topological quantity. The ambiguity in the branch choice of $\alpha_{zz}^{(l)}$ is either fixed by the symmetry constraint on the total bulk response $\sum_{l}\alpha^{(l)}$ or by boundary condition at the surface. We also note that when the total number of layers increases to a certain extent, even a very small electric field applied in the $z$ direction will cause Zener tunnelling and the system then is no longer an insulator. This imposes a truncation to the otherwise diverging layer-resolved OME response.
In summary, we propose a never-mentioned, quantized layer-resolved OME response in generic 3D Chern insulators. This OME response is quantized or half-quantized to $e^{2}/h$ in the presence of inversion or mirror symmetry. Notably, even in the absence of symmetries, the gradient of the OME coupling is still exactly quantized in unit of $e^{2}/h$ given by Eq. (5). We demonstrate the robustness of the quantization rule against disorder through numerical calculations and topological argument using the Streda formula. Such considerable and quantized difference between layer-resolved OME $\alpha_{zz}^{(l)}$ is in principle measurable in experiments. For example, in a slab system, applying an out-of-plane electric field would induce a quantized change in orbital magnetization at the top and bottom layers, which may be probed by SQUID-on-tip microscope [40, 41].
@@ Acknowledgements.
This work is supported by the National Natural Science Foundation of China (grant No. 12174257) and the National Key R & D program of China (grant No. 2020YFA0309601).
Supplemental Materials of “Orbital magnetoelectric coupling of three dimensional Chern Insulator”
@@ I.2 Definition of layer-resolved orbital magnetization in the thermodynamic limit
$-\frac{e}{2\hbar A}{\rm Im}\big{\{}\epsilon_{z\alpha\beta}\sum_{s\,l}\sum_{\textbf{k}}\sum_{n\in{\rm occ}}\sum_{m,m^{\prime}\in{\rm emp}}C_{sl,m}(\textbf{k})C_{sl,m^{\prime}}^{*}(\textbf{k})E_{n\textbf{k}}\langle u_{m\textbf{k}}|\partial_{k_{\alpha}}u_{n\textbf{k}}\rangle_{\Omega_{0}}\langle\partial_{k_{\beta}}u_{n\textbf{k}}|u_{m^{\prime}\textbf{k}}\rangle_{\Omega_{0}}\big{\}}$
@@ I.3 Layer Chern number
In this section, we show that $M_{z}^{(l)}$ varies linearly with weak electric field for both types of stacking, independent of number of layers. Here, we calculate $M_{z}^{(l)}$ using the definition Eq. (23).
@@ VI.3 Layer Chern number
VII VII. Off-diagonal OME response of Chern insulators
Let us consider a 3D Chern insulator with Chern vector $\textbf{C}=(0,0,C_{z})$, where $C_{z}=C(C\in\mathbb{Z},C\neq 0)$ with interlayer distance $d$. What would happen if a non-disruptive in-plane electric field is applied assuming the interlayer coupling is weak? Since 3D Chern insulator is adiabatically connected to trivial stacking of 2D Chern insulators, the response to in-plane electric field (corresponding to off-diagonal $\alpha_{xz}$ or similar components) should behave similarly as in the 2D case. Specifically, if a uniform electric field $\mathcal{E}_{x}$ is applied along the $x$ axis, a transverse anomalous Hall current flowing along the $y$ direction would be induced due to the non-zero Chern number, with $j_{y}=Ce^{2}/(dh)\mathcal{E}_{x}$. The bulk anomalous Hall current would naturally generate orbital magnetization $\mathbf{M}$ via ×$\nabla\times\mathbf{M}=\mathbf{j}$, giving rise to an off-diagonal OME response, with its response coefficient varying linearly in real space, i.e., $\partial_{x}\alpha_{xz}=-Ce^{2}/hd$. Here $e$ is fundamental charge, $h$ is Planck constant, and $d$ is interlayer distance.